OpenStudy (anonymous):

Use the point-slope form of the linear equation to find an equation of the line with the given slope and passing through the given point. Then write the equation in standard form. Slope -4; through (-7, -5) . a. 4x + y = 33 b. 4x + y = -27 c. 4x + y = -5 d. 4x + y = -33

OpenStudy (anonymous):

I love your videos. Sorry. I have no answer.

OpenStudy (anonymous):

Haha Thank you! But i am only a fan of justine's. :D

hero (hero):

Begin by writing an equation in point slope form: (y - y1) = m(x - x1). Then manipulate it until you have the form y = mx + b. In this particular case, you're given the point (x,y) = (-7,-5) and the slope m = -4. Rewriting this in point slope form produces: (y + 5) = -4( x + 7) . To re-write this equation in slope-intercept form, manipulate the equation as follows: (y + 5) = -4( x + 7) y + 5 = -4x - 28 y = -4x - 28 - 5 y = -4x - 33 The last form y = -4x - 33 is in the slope intercept form y = mx +b.